# Staircase Cascaded Fusion of Lightweight Local Pattern Recognition and Long-Range Dependencies for Structural Crack Segmentation

### v2

@@ III-E Loss Function
$L=\alpha\cdot L_{BCE}+\beta\cdot L_{Dice}$ (7)
where $\alpha$ and $\beta$ are hyperparameters that determine the weights of the two losses. In our network, $\alpha$ is set to 0.75 and $\beta$ to 0.25.
@@ IV Experiments
In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics. We then analyze results and model complexity, perform ablation experiments, and discuss current limitations and considerations.
@@ IV-A Dataset Description
4 Crack500 [22] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen
@@
The TUT dataset includes a substantial number of crack images captured with mobile phones, along with a smaller collection of images sourced from the internet. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
As list in Table I, our collected TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 7 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
@@ IV-B Implementation Details
2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation algorithms and the latest SOTA methods on different datasets. These methods include HED [45], UNet++ [46], Deeplabv3+ [47], AttuNet [48], RCF [49], RIND [50], UCTNet [6], SFIAN [7], CT-crackseg [8], DTrCNet [9], Crackmer [10] and Simcrack [11].
$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (9)
where $N$ represents the total number of images.
IV-C Comparison with the SOTA methods
TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Results are highlighted with the best in red and the second-best in blue
@@ IV-C Comparison with the SOTA methods
4) Results on the Crack500 dataset: Similar to the DeepCrack dataset, the Crack500 dataset also has relatively wide crack areas, but its structure is more complex, posing higher requirements on model performance. As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R. It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively. Although DTrCNet [9] leads on the R metric, our method surpasses it on the P metric by 10.51%. The highest F1 score achieved by our method indicates a good balance between P and R. Additionally, as seen in Fig. 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
@@
IV-D Complexity Analysis
To assess our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [9]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines minimal parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
TABLE III: Comparison of the complexity of CrackSCF with other methods. Results are highlighted with the best in red and the second-best in blue
@@ IV-D Complexity Analysis
TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Results are highlighted with the best in red and the second-best in blue
@@
CrackSCF_Original 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
CrackSCF_DS 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
CrackSCF_LR 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
CrackSCF_LRDS 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE V: Comparison of CrackSCF Networks using SCFM, Concat and PAF. Results are highlighted with the best in red and the second-best in blue
@@
CrackSCF_Concat 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
CrackSCF_PAF 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
CrackSCF_SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE VI: Comparison of results of CrackSCF networks trained with different loss ratios
Dice : BCE ODS OIS P R F1 mIoU
1 : 5 0.8129 0.8203 0.8217 0.8458 0.8336 0.8424
1 : 4 0.8049 0.8145 0.8116 0.8365 0.8239 0.8368
1 : 3 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
1 : 2 0.8146 0.8211 0.8227 0.8435 0.8330 0.8438
1 : 1 0.8136 0.8204 0.8231 0.8436 0.8332 0.8431
Only BCE 0.8003 0.8056 0.8349 0.8004 0.8173 0.8332
Only Dice 0.8107 0.8165 0.8302 0.8265 0.8283 0.8409
IV-E Ablation Analysis
@@
To evaluate the advantages of our LRDS block compared to other convolution operations, we performed ablation experiments on the TUT dataset, as list in Table IV. On six metrics, our method slightly outperforms those using only depthwise separable or low-rank approximation convolutions. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the low-rank approximation block, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that our convolution block effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
@@
To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V. Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics. Using only the PAF branch results in a 1.06% drop in the F1 score. While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance. These results indicate that our module effectively balances performance and efficiency.
Table VI shows the results of ablation experiments on different weight proportions of the two sub-losses in the loss function. As shown, a Dice loss to BCE loss ratio of 1:3 yields the highest values in five metrics: ODS, OIS, R, F1, and mIoU. Although the P metric is 0.80% lower than using only BCE loss, the F1 metric improves by 2.58%, indicating a better balance between P and R. Therefore, we use this ratio during training. In summary, a Dice loss to BCE loss ratio of 1:3 not only enhances pixel-level performance but also improves the global segmentation of crack areas, achieving superior segmentation performance.
IV-F Limitations and Discussions
Through these experiments, we found that SCFM effectively perceives and fuses local details and long-range pixel dependencies. Our network excels in various metrics on public datasets with simple scenes and performs exceptionally on the complex TUT dataset. The lightweight LRDS block minimizes parameters and computational load, facilitating deployment on resource-constrained devices like drones and smartphones. However, our network has several limitations, as follows:
  1. 1.
    Although our method generally achieves good crack segmentation with effective background noise suppression, as shown in the Fig. 10, our model occasionally misdetects noise regions that resemble cracks in some images. It also fails to detect watermark occlusions with colors similar to the background. Therefore, it is essential to enhance modules such as MFE and LDE to improve the model’s capability to perceive and extract features from critical noise and occlusion areas, thereby further increasing its robustness and stability in complex scenarios.
  2. 2.
    Despite our method achieves the minimum parameter count and computational load, there is still room for improvement in inference speed. For example, our network achieves a processing speed of 36 frames per second, but the DTrCNet [9], which focuses more on inference speed, attains a faster speed of 47 frames per second. Therefore, it is necessary to further simplify convolution operations or the network’s complexity to achieve faster inference speeds while maintaining a good balance between segmentation performance and model lightweighting.
Figure 10: Examples of CrackSCF misdetections and misses. The critical regions are marked with blue boxes.
@@ V Conclusion
This paper proposes a staircase cascaded fusion network (CrackSCF) for pixel-level crack detection. We developed SCFM to fuse the local patterns processed by MFE and the pixel dependencies processed by LDE. Additionally, we introduced the lightweight convolution block LRDS to replace all convolution operations in the network, effectively reducing the substantial parameter count and computational load required for training and inference. Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes. Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively, with an inference speed of up to 36 frames per second.
In future work, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results. Additionally, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications.

### v3

@@ III-E Loss Function
$L=0.75\cdot L_{B}+0.25\cdot L_{D}$ (15)
where $\mathit{L_{B}}$ represents the BCE loss function and $\mathit{L_{D}}$ represents the Dice loss function.
@@ IV Experiments
In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics. We then analyze results and model complexity, perform ablation experiments.
@@ IV-A Dataset Description
4 Crack500 [23] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen, cement
@@
The TUT dataset comprises 1270 crack images captured with our mobile phones, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex, noisy backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
As list in Table I, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 7 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
@@ IV-B Implementation Details
2) Comparison Methods: In evaluating model performance, we compared the CrackSCF network with 12 classic segmentation methods and the latest SOTA methods on different datasets. These methods include HED [44], UNet++ [45], Deeplabv3+ [46], AttuNet [47], RCF [48], RIND [49], UCTNet [5], SFIAN [6], CT-crackseg [7], DTrCNet [8], Crackmer [9] and Simcrack [10].
TABLE II: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
@@
$ODS=\max_{t}\frac{2\cdot P_{t}\cdot R_{t}}{P_{t}+R_{t}}$ (17)
where $N$ represents the total number of images.
IV-C Comparison with the SOTA methods
@@
4) Results on the Crack500 dataset: Similar to the DeepCrack, the Crack500 dataset also has relatively wide crack areas, but its structure is complex, posing higher requirements on model performance. As list in Table II, our method achieved the best results on five metrics for the Crack500 dataset, except for R. It outperformed the second-best methods on ODS, OIS, P, F1, and mIoU by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%, respectively. Although DTrCNet [8] leads on the R metric, our method surpasses it on the P metric by 10.51%. The highest F1 score achieved by our method indicates a good balance between P and R. Additionally, as seen in Fig. 8, our method does not produce false detections in irrelevant background areas, whereas other methods either produce false detections or lack detailed segmentation in crack areas.
@@
TABLE III: Comparison of the complexity of CrackSCF with other methods. Best results are bolded and second best results are underlined
@@
TABLE IV: Comparison of CrackSCF using LRDS blocks with networks using other convolutions. Best results are bolded and second best results are underlined
@@
Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE V: Comparison of CrackSCF Networks using SCFM, signle Concat branch and signle PAF branch. Best results are bolded and second best results are underlined
@@
Signle Concat branch 0.8019 0.8092 0.8228 0.8209 0.8219 0.8359 7.21G 4.71M 42
Signle PAF branch 0.8137 0.8202 0.8282 0.8297 0.8293 0.8427 8.24G 4.72M 39
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE VI: Comparison of SCFM performance with other segmentation heads. Best results are bolded and second best results are underlined
SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777 36.69G 5.44M 21
SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983 29.72G 6.20M 23
UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222 40.59G 12.34M 20
Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969 168.23G 7.45M 14
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
TABLE VII: Comparison of the performance of using our lightweight deformable attention in LDE with other variants. Best results are bolded and second best results are underlined
Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279 15.98G 5.80M 16
MHA + LR_Linear 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286 8.87G 4.77M 14
Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373 16.37G 5.82M 28
MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
IV-D Complexity Analysis
To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table III presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs (9.26G) and Params (4.79M), surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [8]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
IV-E Ablation Studies
To evaluate the strengths of our proposed module over other modules, we performed ablation experiments on the TUT dataset.
@@ IV-E Ablation Studies
1) Ablation study of different convolution types: As list in Table IV, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
@@
2) Ablation study in SCFM using different branch: To demonstrate the effectiveness of SCFM in fusing local detail features and long-range information, we performed ablation experiments on the TUT dataset, as list in Table V. Our method, using Concat and PAF dual-branch feature fusion, achieves the best results across all six segmentation metrics. Using only the PAF branch results in a 1.06% drop in the F1 score. While SCFM slightly increases parameter count, computational load, and inference speed, it significantly improves segmentation performance. These results indicate that our module effectively balances performance and efficiency.
3) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [50], SegFormer [51], UNet [52], and Hamburger head [53], using the TUT dataset. As list in the Table VI, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
4) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in the Table VII, our method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
@@ V Conclusion
This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection. In CrackSCF, we propose the lightweight convolutional block LRDS to construct the MFE, which dramatically reduces the amount of computation and parametric quantities required by the model and is able to efficiently capture the local details of the crack image, and we propose LDE to more efficiently perceive the pixel irregular long-range dependencies in the crack feature maps, in order to more comprehensively fuse the local patterns with the pixel’s long-range dependencies, we also propose SCFM to fully perceive them to generate higher quality segmentation maps. Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes. Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
Despite our method achieves the low parameter count and computational load, there is still room for improvement in inference speed. In future work, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications. Additionally, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.